OpenStudy (anonymous):

Solve the system of equations: what is the x-value for the solution? 2x + 7y = 3 and x = 1 - 4y

OpenStudy (anonymous):

2(1-4y) + 7y = 3 2 - 8y + 7y = 3 2 + 7y = 3 + 8y 7y = 1 + 8y -1y = 1 y = -1

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

its right thank you:)